# PhySense: Sensor Placement Optimization for Accurate Physics Sensing

### v2

@@ Abstract
Physics sensing plays a central role in many scientific and engineering domains, which inherently involves two coupled tasks: reconstructing dense physical fields from sparse observations and optimizing scattered sensor placements to observe maximum information. While deep learning has made rapid advances in sparse-data reconstruction, existing methods generally omit optimization of sensor placements, leaving the mutual enhancement between reconstruction and placement on the shelf. To change this suboptimal practice, we propose PhySense, a synergistic two-stage framework that learns to jointly reconstruct physical fields and to optimize sensor placements, both aiming for accurate physics sensing. The first stage involves a flow-based generative model enhanced by cross-attention to adaptively fuse sparse observations. Leveraging the reconstruction feedback, the second stage performs sensor placement via projected gradient descent to satisfy spatial constraints. We further prove that the learning objectives of the two stages are consistent with classical variance-minimization principles, providing theoretical guarantees. Extensive experiments across three challenging benchmarks, especially a 3D geometry dataset, indicate PhySense achieves state-of-the-art physics sensing accuracy and discovers informative sensor placements previously unconsidered.
@@ 1 Introduction
Physics sensing remains a foundational task across several domains [44, 22], including fluid dynamics [31, 14], meteorology [38, 4], and industrial applications [12, 3]. The task aims to reconstruct the spatiotemporal information of a physical system from limited sparse observations gathered by pre-placed sensors. In real-world scenarios, the number of sensors is usually limited due to spatial constraints, power consumption, and environmental restrictions. As a result, allocating the limited sensors to the most informative positions is crucial for high reconstruction accuracy. Poorly placed sensors lead to spatial blind spots and information loss, resulting in insufficient observations, while well-placed sensors enable the recovery of fine-scale physical patterns. Therefore, advancing accurate physics sensing requires not only improving the capabilities of reconstruction models, but also developing principled strategies to determine optimal sensor placements.
Recent advances in deep learning have demonstrated remarkable capabilities in sparse-data reconstruction, owing to its ability to approximate highly nonlinear mappings [15, 43, 36] between scattered observations and dense physical fields. Current reconstruction methods fall into two categories: deterministic and generative. Deterministic methods such as VoronoiCNN [13] employ Voronoi-tessellated grids for CNN-based reconstruction, while Senseiver [40] utilizes implicit neural representations [18] to establish correlations between sparse measurements and query points. Moreover, due to the inherently ill-posed nature [21] of the reconstruction task, where sparse observations provide insufficient information for recovering dense fields, generative models [16, 41, 29, 28] have been increasingly explored. For example, DiffusionPDE [17] and $\text{S}^{3}\text{GM}$ [27] model physical field distributions via diffusion models, and integrate sensor observations during sampling through training-free guidance [6, 47]. Despite their reconstruction accuracy, these methods share a fundamental limitation: in their practical experiments, sensor placements are either randomly distributed or arbitrarily fixed rather than being systematically optimized. Consequently, the synergistic potential of co-optimizing reconstruction quality and placement strategy remains largely unexplored.
@@
Building upon these insights, this paper presents PhySense, a synergistic two-stage framework that combines physical field reconstruction with sensor placement optimization. In the first stage, we train a base reconstruction model using flow matching [28] capable of generating dense fields from all feasible sensor placements. In the second stage, leveraging the reconstruction feedback, sensors are optimized using projected gradient descent to satisfy spatial geometries. Furthermore, we prove that the learning objectives of the two stages are consistent with classical variance-minimization principles, providing theoretical guarantees. We evaluate our method across three challenging benchmarks, including turbulent flow simulations, reanalysis of global sea temperature, and industrial simulations of aerodynamic surface pressure over a 3D car on an irregular geometry. In all cases, PhySense consistently achieves state-of-the-art performance and discovers some informative sensor placements previously unconsidered. Overall, our contributions can be summarized as follows:
@@ 2.1 Deep Reconstruction Models
Recent deep learning methods have demonstrated promising performance in reconstructing dense physical fields from sparse measurements. Existing reconstruction models can be broadly categorized into deterministic and generative approaches. Representative deterministic methods are VoronoiCNN [13] and Senseiver [40]. VoronoiCNN constructs a Voronoi tessellation from scattered sensor observations, and learns to map the resulting structured representation to the target physical field. Senseiver encodes arbitrarily sized scattered inputs into a latent space using cross-attention and adopts an implicit neural representation as its decoder. However, since sparse reconstruction is an inherently ill-posed problem, these deterministic methods lack the ability to characterize uncertainty in the reconstructed fields. Deep generative models have also been applied to the sparse reconstruction task. For example, $\text{S}^{3}\text{GM}$ [27] and DiffusionPDE [17] both tackle the reconstruction task by first learning a generative model of the underlying dynamics and then guiding the sampling process using sparse sensor observations. Crucially, these methods only utilize sparse observations during the sampling process—not during training—thus requiring thousands of sampling steps to align with sparse inputs, which compromises the efficiency of deep models. Despite promising reconstruction accuracy, all their experiments are conducted under random or fixed sensor placements, which leads to suboptimal reconstruction accuracy due to potentially non-informative sensor placements.
@@ 2.2 Sensor Placement Optimization
Optimal sensor placement aims to maximize information extraction from physical systems by strategically positioning sensors. This problem is rooted in optimal experimental design (OED) theory [32, 35], which provides basic criteria for evaluating sensor placements. Three most widely-used criteria are: (1) A-optimality, which minimizes the average variance of all data estimates; (2) D-optimality, which maximizes the overall information content; and (3) E-optimality, which focuses on the worst-case estimation error. More details about these criteria can be found in [31]. In this work, we primarily focus on A-optimality. While these criteria originated in classical OED, they have been adapted to sensor placement optimization by incorporating spatial constraints [25].
@@
Sensor placement optimization is a fundamentally hard problem due to its inherent high-dimensional nature. For finite option problems, an exhaustive search over all feasible placements may be tractable, but for more common high-dimensional problems, principled approximations become necessary. The theories of compressed sensing [8] show that signals admitting sparse representations in a known basis, such as the Fourier basis, can be accurately reconstructed with a small number of sensor observations. However, in real-world applications, the assumption of a suitable and known basis may not hold. Moreover, recent advances in machine learning have enabled data-driven discovery of low-dimensional structures such as low-rank subspaces [10, 45, 26]. This shift from fixed to data-driven representation spaces has spurred the development of placement optimization methods that leverage empirical priors or learned latent bases for efficient compression and reconstruction [31, 39]. For example, the SSPOR algorithm [31] first learns a data-driven basis, then applies QR factorization to select basis-specific sensor locations that maximize reconstruction performance. While recent work has begun to explore statistical-based sensor placement, deep learning-based approaches remain under explored. Besides, the further challenge lies in how to effectively co-optimize the reconstruction and the placement for accurate physics sensing.
@@ 2.3 Flow-based Models
Flow matching [28, 29, 30, 1] is a framework for solving transport mapping problem: given two distributions $\mathbf{X}_{0}\sim\pi_{0}$ and $\mathbf{X}_{1}\sim\pi_{1}$, the goal is to find a continuous mapping $T$, such that $T(\mathbf{X}_{0})\sim\pi_{1}$. The method’s efficiency and capacity have made it particularly valuable in vision tasks [11]. Unlike diffusion models that require computationally intensive ODE [41] or SDE [16, 42, 23] solving, flow matching introduces an ODE model that transfers $\pi_{0}$ to $\pi_{1}$ optimally via the straight line, theoretically the shortest path between two points, $\mathrm{d}\mathbf{X}_{t}=\mathbf{v}_{\theta}(\mathbf{X}_{t},t)\,\mathrm{d}t,t\in[0,1],$ where $\mathbf{v}_{\theta}(\mathbf{X}_{t},t)$ is a learnable velocity field. To estimate $\mathbf{v}$, flow matching solves a simple least squares regression problem that fits $\mathbf{v}_{\theta}$ to $(\mathbf{X}_{1}-\mathbf{X}_{0})$. The resulting dynamics generate samples via efficient push-forward along the theoretically shortest path, $\mathbf{X}_{t+\mathrm{d}t}=\mathbf{X}_{t}+\mathbf{v}_{\theta}(\mathbf{X}_{t},t)\mathrm{d}t$, where $\mathbf{X}_{0}\sim\pi_{0}$. In this paper, we consider to integrate this advanced technique for a more efficient reconstruction model.
@@ 3 Method
Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a bounded domain, where exists $\mathcal{C}$ physical fields of interest. We consider =$\mathcal{M}=\{p_{1},\dots,p_{m}\}$ sensors distributed across $\mathcal{D}$, where each sensor measures the target quantities at its position. Our physics sensing task is to accurately reconstruct the entire physical field $\mathbf{X}_{1}$ based on sparse sensor observations $\mathbf{X}_{1}(\mathbf{p^{\ast}})$ under the optimal sensor placement $\mathbf{p^{\ast}}$.
@@ 3.1 Flow-based Reconstructor
where $\mathbf{v}_{\theta}$ represents our learned velocity model, and $\mathbf{p}$ denotes a sensor placement randomly sampled from a uniform distribution on $\mathcal{D}$. This model design must address two practical challenges: (1) handling diverse physical domains spanning regular grids to irregular meshes, and (2) effectively incorporating sparse observations for consistent reconstruction. For the first challenge, we leverage the state-of-the-art DiT [34, 9] for regular grids and Transolver [46], the most advanced approach for general geometry modeling, for irregular geometries, ensuring optimal parameterization of the velocity field across various domains. The second challenge is resolved through a cross-attention mechanism that dynamically associates scattered sensor measurements with their relevant spatial influence regions during training. This not only provides some physical insights into sensor-field interactions but also enhances generalization across varying number and placements of sensors.
@@ 3.2 Sensor Placement Optimization
Based on the well-trained reconstructor, we formalize placement optimization as a constrained optimization problem efficiently solved through a newly proposed projected gradient descent strategy to respect spatial constraints, yielding near A-optimal [32] placements with theoretical guarantees.
Figure 3: Placement optimization stage: Based on a well-trained reconstructor, sensors are optimized by geometry-constrained projected gradient descent to minimize the flow loss with theoretical guarantees.
@@ 4 Experiments
As presented in Table 1, our experiments encompass both regular grids and unstructured meshes in the 2D and 3D domains. The benchmarks include: (1) Turbulent Flow: numerical simulations of channel flow turbulence following the setup in Senseiver [40]; (2) Sea Temperature: 27-year daily sea surface temperature at $1^{\circ}$ resolution with intricate land constraints from the GLORYS12 reanalysis dataset [20]; and (3) Car Aerodynamics: high-fidelity OpenFOAM simulations [19] of surface pressure over a complex car geometry from ShapeNet [5], where we reconstruct the irregular pressure fields from scattered sensor observations.
@@
We evaluate PhySense via two approaches: (1) comparing the reconstruction model under randomly sampled sensor placements, and (2) comparing sensor placements with frozen PhySense’s reconstruction model. The reconstruction benchmarks include five representative baselines: classical statistical-based method SSPOR [31], deterministic models VoronoiCNN [13] and Senseiver [40], and generative approaches $\text{S}^{3}\text{GM}$ [27] and DiffusionPDE [17]. For placement evaluation, we compare our optimized placement against random sampling and SSPOR’s optimized placement [31].
@@
All methods of each benchmark maintain identical training configurations including epochs, batch sizes, ADAM optimizer [24], and maximum learning rate, while preserving other details as specified in their original papers. For the reconstruction model, we employ DiT [34] for turbulent flow and sea temperature benchmarks, while using Transolver [46] as a Transformer backbone to fit irregular mesh for car aerodynamics benchmark. The placement optimization stage runs for 5 epochs with the frozen reconstruction model. Both two stages employ relative L2 as the metrics for the reconstructed fields. Additional implementation details are provided in Appendix C.
@@ 4.1 Turbulent Flow
The benchmark focuses on reconstructing turbulent flow within a channel using a 2-D slice discretized on a $128\times 48$ regular grid. This dataset is widely used, with classical models such as Senseiver [40] and VoroniCNN [13] serving as benchmarks. We replicate Senseiver’s setup; specifically, 25–300 sensors are randomly selected to train the reconstructor, with sensor positions uniformly sampled across the entire domain. A setting with 30 sensors represents low coverage, while 200 sensors corresponds to high coverage. The model is trained using all available simulation data and evaluated on its ability to reconstruct the corresponding fields under varying sensor placements. Besides, due to the computational cost of DiffusionPDE and $\text{S}^{3}\text{GM}$—each reconstruction involving thousands of model evaluations—we restrict their evaluation to 500 randomly sampled cases.
@@
SSPOR [31] SSPOR 0.3018 0.4429 0.5921 0.6637 0.0756 0.0752 0.0789 0.0719
VoronoiCNN [13] Random 0.3588 0.4870 0.6919 0.7992 0.1496 0.1537 0.2433 0.2817
Senseiver [40] 0.1842 0.2316 0.3740 0.5746 0.0715 0.0732 0.0769 0.0784
$\text{S}^{3}\text{GM}$ [27] 0.1856 0.2016 0.2681 0.6421 0.1115 0.1137 0.1581 0.1795
DiffusionPDE [17] 0.1993 0.3312 0.6604 0.9163 0.0664 0.0775 0.0859 0.1139

### v3

@@ Abstract
Physics sensing plays a central role in many scientific and engineering domains, which inherently involves two coupled tasks: reconstructing dense physical fields from sparse observations and optimizing scattered sensor placements to observe maximum information. While deep learning has made rapid advances in sparse-data reconstruction, existing methods generally omit optimization of sensor placements, leaving the mutual enhancement between reconstruction and placement on the shelf. To change this suboptimal practice, we propose PhySense, a synergistic two-stage framework that learns to jointly reconstruct physical fields and to optimize sensor placements, both aiming for accurate physics sensing. The first stage involves a flow-based generative model enhanced by cross-attention to adaptively fuse sparse observations. Leveraging the reconstruction feedback, the second stage performs sensor placement via projected gradient descent to satisfy spatial constraints. We further prove that the learning objectives of the two stages are consistent with classical variance-minimization principles, providing theoretical guarantees. Extensive experiments across three challenging benchmarks, especially a 3D geometry dataset, indicate PhySense achieves state-of-the-art physics sensing accuracy and discovers informative sensor placements previously unconsidered. Code is available at this repository: https://github.com/thuml/PhySense.
@@ 1 Introduction
Physics sensing remains a foundational task across several domains wang2023scientific ; karniadakis2021physics ; wu2024ropinn , including fluid dynamics manohar2018data ; gu2007iterative ; zhouunisolver , meteorology rouet2019continuous ; carrassi2018data , and industrial applications fortuna2007soft ; brunton2021data . The task aims to reconstruct the spatiotemporal information of a physical system from limited sparse observations gathered by pre-placed sensors. In real-world scenarios, the number of sensors is usually limited due to spatial constraints, power consumption, and environmental restrictions. As a result, allocating the limited sensors to the most informative positions is crucial for high reconstruction accuracy. Poorly placed sensors lead to spatial blind spots and information loss, resulting in insufficient observations, while well-placed sensors enable the recovery of fine-scale physical patterns. Therefore, advancing accurate physics sensing requires not only improving the capabilities of reconstruction models, but also developing principled strategies to determine optimal sensor placements.
Recent advances in deep learning have demonstrated remarkable capabilities in sparse-data reconstruction, owing to its ability to approximate highly nonlinear mappings he2016deep ; vaswani2017attention ; radford2018improving between scattered observations and dense physical fields. Current reconstruction methods fall into two categories: deterministic and generative. Deterministic methods such as VoronoiCNN fukami2021global employ Voronoi-tessellated grids for CNN-based reconstruction, while Senseiver santos2023development utilizes implicit neural representations jaegle2021perceiver to establish correlations between sparse measurements and query points. Moreover, due to the inherently ill-posed nature kabanikhin2011inverse of the reconstruction task, where sparse observations provide insufficient information for recovering dense fields, generative models ho2020denoising ; song2020denoising ; liu2022flow ; lipman2022flow have been increasingly explored. For example, DiffusionPDE huang2024diffusionpde and $\text{S}^{3}\text{GM}$ li2024learning model physical field distributions via diffusion models, and integrate sensor observations during sampling through training-free guidance chung2023diffusion ; ye2024tfg . Despite their reconstruction accuracy, these methods share a fundamental limitation: in their practical experiments, sensor placements are either randomly distributed or arbitrarily fixed rather than being systematically optimized. Consequently, the synergistic potential of co-optimizing reconstruction quality and placement strategy remains largely unexplored.
@@
Building upon these insights, this paper presents PhySense, a synergistic two-stage framework that combines physical field reconstruction with sensor placement optimization. In the first stage, we train a base reconstruction model using flow matching lipman2022flow capable of generating dense fields from all feasible sensor placements. In the second stage, leveraging the reconstruction feedback, sensors are optimized using projected gradient descent to satisfy spatial geometries. Furthermore, we prove that the learning objectives of the two stages are consistent with classical variance-minimization principles, providing theoretical guarantees. We evaluate our method across three challenging benchmarks, including turbulent flow simulations, reanalysis of global sea temperature, and industrial simulations of aerodynamic surface pressure over a 3D car on an irregular geometry. In all cases, PhySense consistently achieves state-of-the-art performance and discovers some informative sensor placements previously unconsidered. Overall, our contributions can be summarized as follows:
@@ 2.1 Deep Reconstruction Models
Recent deep learning methods have demonstrated promising performance in reconstructing dense physical fields from sparse measurements. Existing reconstruction models can be broadly categorized into deterministic and generative approaches. Representative deterministic methods are VoronoiCNN fukami2021global and Senseiver santos2023development . VoronoiCNN constructs a Voronoi tessellation from scattered sensor observations, and learns to map the resulting structured representation to the target physical field. Senseiver encodes arbitrarily sized scattered inputs into a latent space using cross-attention and adopts an implicit neural representation as its decoder. However, since sparse reconstruction is an inherently ill-posed problem, these deterministic methods lack the ability to characterize uncertainty in the reconstructed fields. Deep generative models have also been applied to the sparse reconstruction task. For example, $\text{S}^{3}\text{GM}$ li2024learning and DiffusionPDE huang2024diffusionpde both tackle the reconstruction task by first learning a generative model of the underlying dynamics and then guiding the sampling process using sparse sensor observations. Crucially, these methods only utilize sparse observations during the sampling process—not during training—thus requiring thousands of sampling steps to align with sparse inputs, which compromises the efficiency of deep models. Despite promising reconstruction accuracy, all their experiments are conducted under random or fixed sensor placements, which leads to suboptimal reconstruction accuracy due to potentially non-informative sensor placements.
@@ 2.2 Sensor Placement Optimization
Optimal sensor placement aims to maximize information extraction from physical systems by strategically positioning sensors. This problem is rooted in optimal experimental design (OED) theory principlesexperimentaldesign ; pukelsheim2006optimal , which provides basic criteria for evaluating sensor placements. Three most widely-used criteria are: (1) A-optimality, which minimizes the average variance of all data estimates; (2) D-optimality, which maximizes the overall information content; and (3) E-optimality, which focuses on the worst-case estimation error. More details about these criteria can be found in manohar2018data . In this work, we primarily focus on A-optimality. While these criteria originated in classical OED, they have been adapted to sensor placement optimization by incorporating spatial constraints krause2008near .
@@
Sensor placement optimization is a fundamentally hard problem due to its inherent high-dimensional nature. For finite option problems, an exhaustive search over all feasible placements may be tractable, but for more common high-dimensional problems, principled approximations become necessary. The theories of compressed sensing donoho2006compressed show that signals admitting sparse representations in a known basis, such as the Fourier basis, can be accurately reconstructed with a small number of sensor observations. However, in real-world applications, the assumption of a suitable and known basis may not hold. Moreover, recent advances in machine learning have enabled data-driven discovery of low-dimensional structures such as low-rank subspaces drmac2016new ; willcox2006unsteady ; kutz2016dynamic . This shift from fixed to data-driven representation spaces has spurred the development of placement optimization methods that leverage empirical priors or learned latent bases for efficient compression and reconstruction manohar2018data ; rudy2017data . For example, the SSPOR algorithm manohar2018data first learns a data-driven basis, then applies QR factorization to select basis-specific sensor locations that maximize reconstruction performance. While recent work has begun to explore statistical-based sensor placement, deep learning-based approaches remain under explored. Besides, the further challenge lies in how to effectively co-optimize the reconstruction and the placement for accurate physics sensing.
@@ 2.3 Flow-based Models
Flow matching lipman2022flow ; liu2022flow ; ma2024sit ; albergo2023building is a framework for solving transport mapping problem: given two distributions $\mathbf{X}_{0}\sim\pi_{0}$ and $\mathbf{X}_{1}\sim\pi_{1}$, the goal is to find a continuous mapping $T$, such that $T(\mathbf{X}_{0})\sim\pi_{1}$. The method’s efficiency and capacity have made it particularly valuable in vision tasks esser2024scaling . Unlike diffusion models that require computationally intensive ODE song2020denoising or SDE ho2020denoising ; song2020score ; karras2022elucidating solving, flow matching introduces an ODE model that transfers $\pi_{0}$ to $\pi_{1}$ optimally via the straight line, theoretically the shortest path between two points, $\mathrm{d}\mathbf{X}_{t}=\mathbf{v}_{\theta}(\mathbf{X}_{t},t)\,\mathrm{d}t,t\in[0,1],$ where $\mathbf{v}_{\theta}(\mathbf{X}_{t},t)$ is a learnable velocity field. To estimate $\mathbf{v}$, flow matching solves a simple least squares regression problem that fits $\mathbf{v}_{\theta}$ to $(\mathbf{X}_{1}-\mathbf{X}_{0})$. The resulting dynamics generate samples via efficient push-forward along the theoretically shortest path, $\mathbf{X}_{t+\mathrm{d}t}=\mathbf{X}_{t}+\mathbf{v}_{\theta}(\mathbf{X}_{t},t)\mathrm{d}t$, where $\mathbf{X}_{0}\sim\pi_{0}$. In this paper, we consider to integrate this advanced technique for a more efficient reconstruction model.
@@ 3 Method
Let $\mathcal{D}\subset\mathbb{R}^{d}$ be a bounded domain, where exists $\mathcal{C}$ physical fields of interest. We consider $\mathcal{M}=\{p_{1},\dots,p_{m}\}$ sensors distributed across $\mathcal{D}$, where each sensor measures the target quantities at its position. Our physics sensing task is to accurately reconstruct the entire physical field $\mathbf{X}_{1}$ based on sparse sensor observations $\mathbf{X}_{1}(\mathbf{p^{\ast}})$ under the optimal sensor placement $\mathbf{p^{\ast}}$.
@@ 3.1 Flow-based Reconstructor
where $\mathbf{v}_{\theta}$ represents our learned velocity model, and $\mathbf{p}$ denotes a sensor placement randomly sampled from a uniform distribution on $\mathcal{D}$. This model design must address two practical challenges: (1) handling diverse physical domains spanning regular grids to irregular meshes, and (2) effectively incorporating sparse observations for consistent reconstruction. For the first challenge, we leverage the state-of-the-art DiT peebles2023dit ; dosovitskiy2020image for regular grids and Transolver wu2024transolver , the most advanced approach for general geometry modeling, for irregular geometries, ensuring optimal parameterization of the velocity field across various domains. The second challenge is resolved through a cross-attention mechanism that dynamically associates scattered sensor measurements with their relevant spatial influence regions during training. This not only provides some physical insights into sensor-field interactions but also enhances generalization across varying number and placements of sensors.
@@ 3.2 Sensor Placement Optimization
Based on the well-trained reconstructor, we formalize placement optimization as a constrained optimization problem efficiently solved through a newly proposed projected gradient descent strategy to respect spatial constraints, yielding near A-optimal principlesexperimentaldesign placements with theoretical guarantees.
Figure 3: Placement optimization stage: Based on a well-trained reconstructor, sensors are optimized by geometry‑constrained projected gradient descent to minimize the flow loss with theoretical guarantees.
@@ 4 Experiments
As presented in Table 1, our experiments encompass both regular grids and unstructured meshes in the 2D and 3D domains. The benchmarks include: (1) Turbulent Flow: numerical simulations of channel flow turbulence following the setup in Senseiver santos2023development ; (2) Sea Temperature: 27-year daily sea surface temperature at $1^{\circ}$ resolution with intricate land constraints from the GLORYS12 reanalysis dataset jean2021copernicus ; and (3) Car Aerodynamics: high-fidelity OpenFOAM simulations jasak2009openfoam of surface pressure over a complex car geometry from ShapeNet chang2015shapenet , where we reconstruct the irregular pressure fields from scattered sensor observations.
@@
We evaluate PhySense via two approaches: (1) comparing the reconstruction model under randomly sampled sensor placements, and (2) comparing sensor placements with frozen PhySense’s reconstruction model. The reconstruction benchmarks include five representative baselines: classical statistical-based method SSPOR manohar2018data , deterministic models VoronoiCNN fukami2021global and Senseiver santos2023development , and generative approaches $\text{S}^{3}\text{GM}$ li2024learning and DiffusionPDE huang2024diffusionpde . For placement evaluation, we compare our optimized placement against random sampling and SSPOR’s optimized placement manohar2018data .
@@
All methods of each benchmark maintain identical training configurations including epochs, batch sizes, ADAM optimizer adam , and maximum learning rate, while preserving other details as specified in their original papers. For the reconstruction model, we employ DiT peebles2023dit for turbulent flow and sea temperature benchmarks, while using Transolver wu2024transolver as a Transformer backbone to fit irregular mesh for car aerodynamics benchmark. The placement optimization stage runs for 5 epochs with the frozen reconstruction model. Both two stages employ relative L2 as the metrics for the reconstructed fields. Additional implementation details are provided in Appendix C.
@@ 4.1 Turbulent Flow
The benchmark focuses on reconstructing turbulent flow within a channel using a 2‑D slice discretized on a $128\times 48$ regular grid. This dataset is widely used, with classical models such as Senseiver santos2023development and VoroniCNN fukami2021global serving as benchmarks. We replicate Senseiver’s setup; specifically, 25–300 sensors are randomly selected to train the reconstructor, with sensor positions uniformly sampled across the entire domain. A setting with 30 sensors represents low coverage, while 200 sensors corresponds to high coverage. The model is trained using all available simulation data and evaluated on its ability to reconstruct the corresponding fields under varying sensor placements. Besides, due to the computational cost of DiffusionPDE and $\text{S}^{3}\text{GM}$—each reconstruction involving thousands of model evaluations—we restrict their evaluation to 500 randomly sampled cases.
@@
SSPOR manohar2018data SSPOR 0.3018 0.4429 0.5921 0.6637 0.0756 0.0752 0.0789 0.0719
VoronoiCNN fukami2021global Random 0.3588 0.4870 0.6919 0.7992 0.1496 0.1537 0.2433 0.2817
Senseiver santos2023development 0.1842 0.2316 0.3740 0.5746 0.0715 0.0732 0.0769 0.0784
$\text{S}^{3}\text{GM}$ li2024learning 0.1856 0.2016 0.2681 0.6421 0.1115 0.1137 0.1581 0.1795
DiffusionPDE huang2024diffusionpde 0.1993 0.3312 0.6604 0.9163 0.0664 0.0775 0.0859 0.1139